A geostationary satellite is orbiting the earth at a height of 6 R above the surface of the earth, where R is the radius of the earth. The time period of another satellite at a height of  2.5 R from the surface of the earth is ...........hours ?

Grade:6

1 Answers

Deepak Patra
9 years ago

KEY CONCEPT : According to Kepler’s law T2 ? R3

T21 / T22= R31/ R32 . Here R = R + 6R = 7R

and R2 = 2.5 R + R = 3.5 R

? 24x 24 /T22 = 7 x 7 x 7 R3 / 3.5 x 3.5 x 3.5 x R3 ? T2 = 8.48 hr